QUESTION 6 The mean IQ score of all adults is 100, with a standard deviation of 15. Find the percentage of adults with scores between 70 and 130. Assume the data has a bell-shaped distribution. State the answer only and no additional work. Do not state the percent sign and state as a whole number. (Example: do not state 13% or 0.13. Instead state the answer as 13)
QUESTION 4 2 po Many firms use on-the-job training to teach their employees computer programming. Suppose you work in the personnel department of a firm that just finished training a group of its employees to program, and you have been requested to review the performance of one of the trainees on the final test that was given to all trainees. The mean test score for all trainees is 72 with a standard deviation of 5. Assume the distribution of scores is bell-shaped. Suppose the trainee in question received a score of 65. Computer the trainees z-score. State the answer only and no additional work. Round the answer to the tenths place.
Answer #1

Question 6

Solution:

We are given that the data has a bell shaped (normal) distribution.

Mean = 100

SD = 15

We have to find P(70<X<130)

P(70<X<130) = P(X<130) – P(X<70)

Z = (X – mean) / SD

First find P(X<130)

Z = (130 – 100) / 15

Z = 30/15 = 2

P(Z<2) = P(X<130) = 0.97725

(By using z-table)

Now, find P(X<70)

Z = (70 – 100)/15

Z = -2

P(Z<-2) = P(X<70) = 0.02275

(By using z-table)

P(70<X<130) = P(X<130) – P(X<70)

P(70<X<130) = 0.97725 – 0.02275

P(70<X<130) = 0.9545

Answer: 95

Question 4

We are given

Data follows bell shaped distribution or normal distribution.

Mean = 72

SD = 5

X = 65

Z = (X – mean) / SD

Z = (65 – 72) / 5

Z = -7/5

Z = -1.4

Answer: -1.4